OpenStudy (anonymous):

write in standard form 3y=9x-12

OpenStudy (anonymous):

Standard form is Ax + By = C

OpenStudy (anonymous):

So you move the 9x to the other side to give you 3y-9x=-12

OpenStudy (anonymous):

yes. Now take it a step further and put the X in the front. Not sure of what your teacher requires but it is not good math protocol to leave a negative in the front. So you would divide by negative 1 (each term) in order to make it be proper like

OpenStudy (anonymous):

Standard Form is Ax + By +C =0, or Ax + BY = C For 3y = 9x -12 Divide by 3 y = 3x - 4 3x - y -4 =0 or 3x-y = 4